OpenStudy (anonymous):

A car travels 1/6 of the distance between two cities in 3/5 of an hour. At this rate, what fraction of the distance between the two cities can the car travel in 1 hour? Enter your fraction in the space provided. /

OpenStudy (anonymous):

@phi

OpenStudy (phi):

I would use ratios. distance/ time = distance / time

OpenStudy (anonymous):

so then 1/6 / 3/5 = 3/5 / 1/6

OpenStudy (anonymous):

Right @phi?

OpenStudy (phi):

ok on the "left side" on the right side we have "unknown distance" over 1 hour of time so \[ \frac{ \frac{1}{6}}{\frac{3}{5}} = \frac{x}{1}\]

OpenStudy (anonymous):

so then...

OpenStudy (phi):

x is the unknown distance x/1 is just x to simplify. multiply the top and bottom of the "ugly fraction" by 5/3

OpenStudy (anonymous):

x/1 * 5/3?

OpenStudy (anonymous):

5x/3

OpenStudy (phi):

you could "cross multiply" (maybe that is what you are trying to do ?) but we want x by itself.. \[ \frac{ \frac{1}{6}}{\frac{3}{5}} = \frac{x}{1} \\ \frac{ \frac{1}{6}\cdot \frac{5}{3}}{\frac{3}{5}\cdot \frac{5}{3}} = x \]

OpenStudy (phi):

notice that 3/5 * 5/3 is 1 and 1 in the denominator means we can ignore it.

OpenStudy (anonymous):

5/18 / 15/15

OpenStudy (phi):

15/15 = 1 when you divide by 1, you can ignore it , right ?

OpenStudy (anonymous):

yes

OpenStudy (phi):

so that simplifies to 5/18

OpenStudy (anonymous):

so then that's the answer??

OpenStudy (phi):

yes. if we go back to the original ratio, we are saying \[ \frac{ \frac{1}{6}}{\frac{3}{5}} = \frac{\frac{5}{18}}{1} \] in other words: if we go 1/6 of the way in 3/5 of an hour, then we go 5/18 of the way in 1 hour

OpenStudy (anonymous):

ok, thx!!!!!!!!!!!!!!!

OpenStudy (anonymous):

can u help with another one?

OpenStudy (phi):

last one

OpenStudy (anonymous):

ok :D

OpenStudy (anonymous):

:3 77777jeannie77777 how did u like flvs?